# Finetuning Large Language Models for Vulnerability Detection

### v4

@@
Abstract.
This paper presents the results of finetuning large language models (LLMs) for the task of detecting vulnerabilities in source code. We leverage WizardCoder, a recent improvement of the state-of-the-art LLM StarCoder, and adapt it for vulnerability detection through further finetuning. To accelerate training, we modify WizardCoder’s training procedure, also we investigate optimal training regimes. For the imbalanced dataset with many more negative examples than positive, we also explore different techniques to improve classification performance. The finetuned WizardCoder model achieves improvement in ROC AUC and F1 measures on balanced and imbalanced vulnerability datasets over CodeBERT-like model, demonstrating the effectiveness of adapting pretrained LLMs for vulnerability detection in source code. The key contributions are finetuning the state-of-the-art code LLM, WizardCoder, increasing its training speed without the performance harm, optimizing the training procedure and regimes, handling class imbalance, and improving performance on difficult vulnerability detection datasets. This demonstrates the potential for transfer learning by finetuning large pretrained language models for specialized source code analysis tasks.
Large language models, vulnerability detection, finetuning, StarCoder, WizardCoder, PEFT, LoRA, defect detection, Java dataset, deep learning, software vulnerabilities, vulnerability prediction, vulnerability classification, cybersecurity, AI
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: The 4th International Conference on Code Quality; 22 June, 2024; Innopolis University, Russiaccs: Security and privacy Software and application securityccs: Computing methodologies Machine learning approaches
1. Introduction
@@
1.1. Motivation
Transformer models pretrained on large code corpus, like CodeBERT (Feng et al., 2020) and ContraBERT (Liu et al., 2023), have shown promising results on many tasks of code understanding. LineVul (Fu and Tantithamthavorn, 2022), a model built on CodeBERT, achieved state-of-the-art results on the vulnerability detection task, surpassing graph network approaches, like Reveal (Chakraborty et al., 2022). Recent LLMs, which have billions of parameters, such as StarCoder (Li et al., 2023) and WizardCoder (Luo et al., 2023), are pretrained on thrillions of tokens. They are 2 orders of magnitude bigger than CodeBERT (CodeBERT has approximately 125 million parameters) and have orders of magnitude more data used for pretraining. They achieve better results on language and code understanding tasks, providing an opportunity for improved transfer learning. The vulnerability detection task can benefit from the increased expressiveness and knowledge contained in state-of-the-art LLMs.
Vulnerability detection is a complex task with many aspects influencing and potentially limiting the solution quality. One of the factors could be limits on the capacity and code understanding ability of CodeBERT-like models, as vulnerability detection requires a deep understanding of many code aspects. Checking this hypothesis is a crucial problem in the field of vulnerability detection.
Furthermore, our previous study (Cheshkov et al., 2023) revealed that simple prompting techiques for the ChatGPT LLM were unable to achieve noticeably better results than random prediction on the vulnerability detection task. This finding suggests that finetuning of LLMs on this task naturally becomes a promising research direction.
The real-world distribution of vulnerable functions is highly class imbalanced, containing far more negative samples than positive ones. This imbalance makes the task inherently more difficult, as standard training objectives, like cross-entropy loss, may lead the model to focus on the dominant negative class. Specialized training techniques, such as focal loss (Lin et al., 2017) and sample weighting, are required to properly emphasize the minority positive examples during finetuning.
@@ 1.1. Motivation
  • We describe our approach for building and adapting the model. Specifically, we describe how we solved the following problems:
@@
  • We describe our main instruments: datasets, models, metrics, as well as measurement procedures;
@@
2. Implementation details
2.1. Challenges
@@
Recently, there has been a rise in the number of open-sourced LLM models, trained on large code corpora (Li et al., 2023; Luo et al., 2023; Zheng et al., 2023; Nijkamp et al., 2023). The task is to choose an optimal model for finetuning, both from the perspective of quality and potential technical difficulties encountered. The goal is to select a model possessing the following properties:
@@
2.2. Choosing an LLM to finetune
Recently, there has been a proliferation of open-sourced LLM models trained on large code corpora (Li et al., 2023; Luo et al., 2023; Zheng et al., 2023; Nijkamp et al., 2023). The task is to select an optimal model for finetuning.
First, we attempted to use the CodeGeeX model (Zheng et al., 2023). Unfortunately, we found that using this model was very difficult from the technical perspective, as the code provided by the authors does not support standard libraries for transformer training and large model adaptation. We were able to run the model on 8 V100 GPUs for inference, but this memory capacity was insufficient for finetuning. The lack of support for common LLM adaptation libraries made it very challenging to overcome the memory limitations, so we decided to discontinue attempts with this model.
Next, we explored two other models, WizardCoder (Luo et al., 2023) and CodeGen (Nijkamp et al., 2023). In contrast to CodeGeeX, these models include support for major transformer training frameworks like DeepSpeed, HuggingFace Transformers, Peft (Mangrulkar et al., 2022), etc. This compatibility with standard libraries significantly eased the application of low-memory adaptation techniques.
@@ 2.2. Choosing an LLM to finetune
In our study, we evaluate three models from the StarCoder family for finetuning: StarCoderBase, StarCoder (Li et al., 2023) (which is StarCoderBase finetuned on the Python dataset), and WizardCoder (Luo et al., 2023) (which is an improved version of StarCoder trained using the eval-instruct technique). We compared the performance of StarCoder and StarCoderBase on the question answering task and found no significant differences between the two models, despite the fact that StarCoder was specifically adapted for Python. Subsequently, we primarily used the WizardCoder model, as it is claimed to be superior to StarCoder. However, our experiments did not reveal any significant differences in performance between these two models. In our future writing, we will use the term “StarCoder” to refer specifically to the WizardCoder variant of the StarCoder family.
@@
2.3. Adapting LLMs for limited computational resources
Finetuning complete multi-billion parameter LLMs is infeasible given our hardware constraints, even if more resources were obtained. To enable effective finetuning, we require techniques to reduce the number of trained parameters and memory needs. The HuggingFace Peft (Mangrulkar et al., 2022) library implements several methods tackling this, including the LORA method (Hu et al., 2022). LORA enables full-model finetuning by decomposing the weight matrices into low-rank approximations, drastically decreasing parameters and memory.
@@ 2.3. Adapting LLMs for limited computational resources
2.4. Adapting LLMs for classification
@@
(1) $\mathcal{L}_{\text{ntp}}=-\sum_{n=1}^{N-1}y_{n+1}\log\mathbb{P}(x_{n+1}|x_{1},x_{2},\ldots,x_{n},\Theta)$
@@
(2) $\mathcal{L}_{\text{class}}=-\log\mathbb{P}(y|x_{1},x_{2},\ldots,x_{N},\Theta)$
@@
2.5. Speeding up the classification training
@@
Table 1. Function length histogram, obtained from parsing several projects
@@
  1. (1)
    Next token prediction. This approach corresponds to using the next token prediction loss (1) during training and prediction. During training, an LLM’s input sequence is packed with the code of training methods and their ground truth labels one after another in the following format:
    Method’s code + YES / NO (binary ground truth label) + ¡EOS¿
    Here ¡EOS¿ denotes the special “end of sequence” token from the vocabulary. We fit as many as possible full training examples into the input sequence length, which is equal to $2048$ for StarCoder-based models and $512$ for CodeBERT-based ones. The rest of the positions are filled with the special padding token. For prediction, the binary classification token is generated according to (1), and its probability is taken for ROC AUC calculation;
  2. (2)
    Binary classification. This regime uses the binary cross entropy classification loss (2) for training and prediction. Specifically, the training scheme for binary classification loss is formalized as follows:
@@
      Calculate the cross-entropy loss for each pair of predicted probability and actual label using (2);
@@
    (3) $\mathcal{L}_{\text{batch}}=\sum_{i=1}^{B}-y_{i}\cdot\log(p_{i})-(1-y_{i})\cdot\log(1-p_{i})$
@@
3. Experimental setup
3.1. Datasets
@@
Our vulnerability dataset is constructed from several open-source vulnerability datasets: CVEfixes (Bhandari et al., 2021), a Manually-Curated Dataset (Ponta et al., 2019) and VCMatch (Wang et al., 2022).
@@
More specifically, extracting functions from the commits fixing only one function mitigates labeling errors that stem from irrelevant code and cleanup changes (Croft et al., 2023).
@@
As a result of these procedures, we obtain a set of vulnerable functions (we call it $P_{1}$) and a set of vulnerable functions with fixed vulnerability (we call it $P_{2}$). We can also augment these sets with a set of functions randomly scratched from random projects, which do not have any vulnerabilities. We call this part $P_{3}$.
@@
3.2. Evaluation metrics
@@
3.3. Pretrained models
In this work, the primary pretrained model architecture employed is WizardCoder (Luo et al., 2023), a transformer-based neural network with a decoder-only architecture. With over 13 billion parameters, WizardCoder was pretrained using a causal language modeling objective on a large collection of GitHub source code, endowing the model with extensive knowledge of natural programming language constructs.
We compare WizardCoder to ContraBERT (Liu et al., 2023), which is the state-of-the-art model in the CodeBERT family. In contrast to vanilla CodeBERT, ContraBERT introduces an enhanced encoder architecture obtained by cooperation of two pretrained CodeBERT encoders. ContraBERT is pretrained using the usual masked language modeling (MLM) along with a number of contrastive pretraining tasks, and can be finetuned for the defect detection task (Liu et al., 2023). ContraBERT can be considered as an enhancement of the CodeBERT model, which showcases more accurate and robust results than CodeBERT for the defect detection task. Hence, ContraBERT is selected for the experiments.
3.4. Full reproducibility package
All necessary code and data for reproducing our experiments are available in our GitHub repository [^1] https://github.com/rmusab/vul-llm-finetune.
4. Research questions
4.1. Comparison to CodeBERT-based models.
@@ 4. Research questions
4.3. Imbalanced classification improvement.
@@
5. Experimental results
5.1. Comparison to CodeBERT-based models
@@
Table 2. ROC AUC and F1 scores for finetuned ContraBERT and WizardCoder models on the $X_{1}$ without $P_{3}$ dataset.
@@
Table 3. ROC AUC and F1 scores for finetuned ContraBERT and WizardCoder models on $X_{1}$ with $P_{3}$ dataset.
5.2. Ablation Study
@@
Table 4. ROC AUC Scores for finetuning with the next token prediction and classification objectives on the imbalanced $X_{1}$ with $P_{3}$ dataset.
@@
Table 5. WizardCoder’s results with limiting the maximum number of functions per batch on the validation part of the $X_{1}$ without $P_{3}$ dataset.
@@
Table 6. WizardCoder’s performance by a loss reduction method on the validation part of the $X_{1}$ without $P_{3}$ dataset.
@@
Table 7. The impact of the WizardCoder’s context size on performance.
@@
5.3. Imbalanced classification improvement

### v5

@@
Abstract
This paper presents the results of finetuning large language models (LLMs) for the task of detecting vulnerabilities in Java source code. We leverage WizardCoder, a recent improvement of the state-of-the-art LLM StarCoder, and adapt it for vulnerability detection through further finetuning. To accelerate training, we modify WizardCoder’s training procedure, also we investigate optimal training regimes. For the imbalanced dataset with many more negative examples than positive, we also explore different techniques to improve classification performance. The finetuned WizardCoder model achieves improvement in ROC AUC and F1 measures on balanced and imbalanced vulnerability datasets over CodeBERT-like model, demonstrating the effectiveness of adapting pretrained LLMs for vulnerability detection in source code. The key contributions are finetuning the state-of-the-art code LLM, WizardCoder, increasing its training speed without the performance harm, optimizing the training procedure and regimes, handling class imbalance, and improving performance on difficult vulnerability detection datasets. This demonstrates the potential for transfer learning by finetuning large pretrained language models for specialized source code analysis tasks.
1 Introduction
@@
Transformer models pretrained on large code corpus, like CodeBERT Feng et al. (2020) and ContraBERT Liu et al. (2023a), have shown promising results on many tasks of code understanding. LineVul Fu & Tantithamthavorn (2022), a model built on CodeBERT, achieved state-of-the-art results on the vulnerability detection task, surpassing graph network approaches, like Reveal Chakraborty et al. (2022). Recent LLMs, which have billions of parameters, such as StarCoder Li et al. (2023) and WizardCoder Luo et al. (2023), are pretrained on trillions of tokens. They are 2 orders of magnitude bigger than CodeBERT (CodeBERT has approximately 125 million parameters) and have orders of magnitude more data used for pretraining. They achieve better results on language and code understanding tasks, providing an opportunity for improved transfer learning. The vulnerability detection task can benefit from the increased expressiveness and knowledge contained in state-of-the-art LLMs.
Vulnerability detection is a complex task with many aspects influencing and potentially limiting the solution quality. The problem of function-level vulnerability detection can be divided into two stages: representing the function in some form (such as execution paths or various graphs) and solving the classification problem for this representation Chakraborty et al. (2022); Zhang et al. (2023). It is noted that each stage can be independently improved, and improving each stage enhances the overall quality of the solution to the problem. Justifications to these observations include: the EPVD model Zhang et al. (2023), where improving the data structure with the same ML algorithm resulted in gains, and the LineVul model Fu & Tantithamthavorn (2022) compared to articles that present the function text to an LSTM. Therefore, it is relevant to improve each of the two stages.
We focus on improving the second part, the ML algorithm for vulnerability detection, without addressing the first stage. Our approach involves two directions: enhancing the model’s capacity and improving the handling of imbalanced data, which represents real-world conditions. One of the limiting factors could be the capacity and code understanding ability of CodeBERT-like models, as vulnerability detection requires a deep understanding of many code aspects. Checking this hypothesis is a crucial problem in the field of vulnerability detection. Hence, we compare our classification of functions, represented as text, with the previous most powerful family of models, the CodeBERT-based models.
Furthermore, the study by Cheshkov et al. (2023) revealed that simple prompting techiques for the ChatGPT LLM were unable to achieve noticeably better results than random prediction on the vulnerability detection task. This finding suggests that finetuning of LLMs on this task naturally becomes a promising research direction.
The real-world distribution of vulnerable functions is highly class imbalanced, containing far more negative samples than positive ones. This imbalance makes the task inherently more difficult, as standard training objectives, like cross-entropy loss, may lead the model to focus on the dominant negative class. Specialized training techniques, such as focal loss Lin et al. (2017) and sample weighting, are required to properly emphasize the minority positive examples during finetuning.
@@
  • We discuss the most notable related work in the field;
  • We describe our approach for building and adapting the model. Specifically, we discuss how we solved the following problems:
@@
  • We detail our main instruments: datasets, models, metrics, as well as measurement procedures.
@@
Appendix A provides additional details about the obtained distribution of CWE types in our dataset.
2 Related work
Recent work by Zhou et al. Zhou et al. (2024) provides a comprehensive review of Large Language Models (LLMs) for vulnerability detection and repair. Their analysis reveals rapid growth in this field, with 27.8% of reviewed studies published in just the first two months of 2024. The majority of approaches (82%) employ fine-tuning techniques, while zero-shot (11%) and few-shot (7%) methods are less common. CodeBERT emerges as the dominant model for vulnerability detection, used in 38.5% of studies, with GPT-3.5 and GPT-4 following at 10.3% and 2.6% respectively. The review also highlights persistent challenges in the field, including class imbalance, noisy data, and scarcity of labeled datasets for training LLMs in vulnerability detection. Our work addresses these challenges by focusing on fine-tuning WizardCoder for Java vulnerability detection, contributing to this rapidly evolving research area.
Despite the successes of GPT models in solving various code-related tasks, their ability to detect vulnerabilities in code remains insufficient. The report ”Evaluation of ChatGPT Model for Vulnerability Detection” (Cheshkov et al., 2023) notes that ChatGPT and GPT-3 models, without additional tuning and training, show results no better than random guessing in vulnerability classification. Similar conclusions are drawn in the study ”ChatGPT for Vulnerability Detection, Classification, and Repair: How Far Are We?” (Michael Fu et al.) (Fu et al., 2023), where comparisons with specialized models such as CodeBERT and GraphCodeBERT reveal a significant lag of ChatGPT in vulnerability-related tasks.
Another notable example is the study by Liu et al. (2023) (Liu et al., 2023b), where the GPT model was enhanced using retrieval-augmented generation (RAG). Techniques such as BM-25 and TF-IDF were employed to retrieve relevant information, which was then used as examples for the few-shot learning technique. Using the Devign dataset, the study demonstrated significant improvements in precision, recall, and F1 score compared to traditional methods.
Recent work by Purba et al. Purba et al. (2023) explores the application of large language models (LLMs) for software vulnerability detection in C and C++ code. They evaluate four LLMs, including fine-tuned and zero-shot models, on SQL injection and buffer overflow vulnerabilities using the Code Gadgets and CVEfixes datasets. Their findings reveal that while LLMs demonstrate high recall in identifying vulnerable code patterns, they suffer from high false positive rates. This aligns with our observations on the challenges of using LLMs for vulnerability detection. However, our study extends this approach to Java, a domain where LLM-based vulnerability detection remains underexplored. We focus specifically on fine-tuning the state-of-the-art WizardCoder model for Java code analysis and introduce novel techniques to improve training efficiency and handle class imbalance. Unlike Purba et al., who primarily use prompt-based approaches, we explore more advanced fine-tuning methods and provide a detailed analysis of the model’s performance on Java-specific vulnerability types, contributing to the limited body of work on LLM-based vulnerability detection for Java.
3 Implementation details
3.1 Challenges
@@
Recently, there has been a rise in the number of open-sourced LLM models, trained on large code corpora Li et al. (2023); Luo et al. (2023); Zheng et al. (2023); Nijkamp et al. (2023). The task is to choose an optimal model for finetuning, both from the perspective of quality and potential technical difficulties encountered. The goal is to select a model possessing the following properties:
@@
3.2 Choosing an LLM to finetune
Recently, there has been a proliferation of open-sourced LLM models trained on large code corpora Li et al. (2023); Luo et al. (2023); Zheng et al. (2023); Nijkamp et al. (2023). The task is to select an optimal model for finetuning.
First, we attempted to use the CodeGeeX model Zheng et al. (2023). Unfortunately, we found that using this model was very difficult from the technical perspective, as the code provided by the authors does not support standard libraries for transformer training and large model adaptation. We were able to run the model on 8 V100 GPUs for inference, but this memory capacity was insufficient for finetuning. The lack of support for common LLM adaptation libraries made it very challenging to overcome the memory limitations, so we decided to discontinue attempts with this model.
Next, we explored two other models, WizardCoder Luo et al. (2023) and CodeGen Nijkamp et al. (2023). In contrast to CodeGeeX, these models include support for major transformer training frameworks like DeepSpeed, HuggingFace Transformers, Peft Mangrulkar et al. (2022), etc. This compatibility with standard libraries significantly eased the application of low-memory adaptation techniques.
@@ 3.2 Choosing an LLM to finetune
In our study, we evaluate three models from the StarCoder family for finetuning: StarCoderBase, StarCoder Li et al. (2023) (which is StarCoderBase finetuned on the Python dataset), and WizardCoder Luo et al. (2023) (which is an improved version of StarCoder trained using the eval-instruct technique). We compared the performance of StarCoder and StarCoderBase on the question answering task and found no significant differences between the two models, despite the fact that StarCoder was specifically adapted for Python. Subsequently, we primarily used the WizardCoder model, as it is claimed to be superior to StarCoder. However, our experiments did not reveal any significant differences in performance between these two models. In our future writing, we will use the term “StarCoder” to refer specifically to the WizardCoder variant of the StarCoder family.
@@
3.3 Adapting LLMs for limited computational resources
Finetuning complete multi-billion parameter LLMs is infeasible given our hardware constraints, even if more resources were obtained. To enable effective finetuning, we require techniques to reduce the number of trained parameters and memory needs. The HuggingFace Peft Mangrulkar et al. (2022) library implements several methods tackling this, including the LORA method Hu et al. (2022). LORA enables full-model finetuning by decomposing the weight matrices into low-rank approximations, drastically decreasing parameters and memory.
@@ 3.3 Adapting LLMs for limited computational resources
3.4 Adapting LLMs for classification
@@
$\mathcal{L}_{\text{ntp}}=-\sum_{n=1}^{N-1}y_{n+1}\log\mathbb{P}(x_{n+1}|x_{1},x_{2},\ldots,x_{n},\Theta)$ (1)
@@
$\mathcal{L}_{\text{class}}=-\log\mathbb{P}(y|x_{1},x_{2},\ldots,x_{N},\Theta)$ (2)
@@
3.5 Speeding up the classification training
@@
Table 1: Function length histogram, obtained from parsing several projects
@@
  1. 1.
    Next token prediction. This approach corresponds to using the next token prediction loss equation 1 during training and prediction. During training, an LLM’s input sequence is packed with the code of training methods and their ground truth labels one after another in the following format:
[…]
  2. 2.
    Binary classification. This regime uses the binary cross entropy classification loss equation 2 for training and prediction. Specifically, the training scheme for binary classification loss is formalized as follows:
@@
      Calculate the cross-entropy loss for each pair of predicted probability and actual label using equation 2;
@@
    $\mathcal{L}_{\text{batch}}=\sum_{i=1}^{B}-y_{i}\cdot\log(p_{i})-(1-y_{i})\cdot\log(1-p_{i})$ (3)
@@
4 Experimental setup
4.1 Datasets
@@
Our vulnerability dataset is constructed from several open-source vulnerability datasets: CVEfixes Bhandari et al. (2021), a Manually-Curated Dataset Ponta et al. (2019) and VCMatch Wang et al. (2022).
@@
More specifically, extracting functions from the commits fixing only one function mitigates labeling errors that stem from irrelevant code and cleanup changes Croft et al. (2023).
@@
As a result of these procedures, we obtain a set of vulnerable functions (we call it $P_{1}$) and a set of vulnerable functions with fixed vulnerability (we call it $P_{2}$). We augmented these sets with a set of unchanged functions scratched from files associated with patch, which do not have any vulnerabilities. We call this part $P_{3}$.
@@
In Appendix A, Figures 1 and 2 show the distribution of the top 19 CWE types for the resulting test datasets with and without the $P_{3}$ part.
4.2 Evaluation metrics
@@
4.3 Pretrained models
In this work, the primary pretrained model architecture employed is WizardCoder Luo et al. (2023), a transformer-based neural network with a decoder-only architecture. With over 13 billion parameters, WizardCoder was pretrained using a causal language modeling objective on a large collection of GitHub source code, endowing the model with extensive knowledge of natural programming language constructs.
We compare WizardCoder to ContraBERT Liu et al. (2023a), which is the state-of-the-art model in the CodeBERT family. In contrast to vanilla CodeBERT, ContraBERT introduces an enhanced encoder architecture obtained by cooperation of two pretrained CodeBERT encoders. ContraBERT is pretrained using the usual masked language modeling (MLM) along with a number of contrastive pretraining tasks, and can be finetuned for the defect detection task Liu et al. (2023a). ContraBERT can be considered as an enhancement of the CodeBERT model, which showcases more accurate and robust results than CodeBERT for the defect detection task. Hence, ContraBERT is selected for the experiments.
4.4 Full reproducibility package
All necessary code and data for reproducing our experiments are available in our GitHub repository [^1] Source code and dataset: https://github.com/rmusab/vul-llm-finetune.
5 Research questions
5.1 Comparison to CodeBERT-based models.
@@ 5 Research questions
5.3 Imbalanced classification improvement.
@@
6 Experimental results
6.1 Comparison to CodeBERT-based models
@@
Table 2: ROC AUC and F1 scores for finetuned ContraBERT and WizardCoder models on the $X_{1}$ without $P_{3}$ dataset.
@@
Table 3: ROC AUC and F1 scores for finetuned ContraBERT and WizardCoder models on $X_{1}$ with $P_{3}$ dataset.
6.2 Ablation study
@@
Table 4: ROC AUC Scores for finetuning with the next token prediction and classification objectives on the imbalanced $X_{1}$ with $P_{3}$ dataset.
@@
Table 5: WizardCoder’s results with limiting the maximum number of functions per batch on the validation part of the $X_{1}$ without $P_{3}$ dataset.
@@
Table 6: WizardCoder’s performance by a loss reduction method on the validation part of the $X_{1}$ without $P_{3}$ dataset.
@@
Table 7: The impact of the WizardCoder’s context size on performance.
@@
6.3 Imbalanced classification improvement